OpenStudy (anonymous):

Mike was working on solving the exponential equation 37^x = 12; however, he is not quite sure where to start. Using complete sentences, describe to Mike how to solve this equation and how solving would be different if the bases were equal. @shamil98 @kc_kennylau

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

this question is so meta

OpenStudy (anonymous):

@Phaen what do you mean

OpenStudy (anonymous):

@Phaen

OpenStudy (anonymous):

you'll want to use logarithms but logs are my weakness >.< I gotta read awhile to make sure I got my facts straight

OpenStudy (anonymous):

:/ thank you for helping I really need help. Tag someone you know that can help

OpenStudy (anonymous):

I'm trying to remember how to rewrite the exponential equation as a logarithmic one

OpenStudy (anonymous):

oh I think I can tell you hold up.

OpenStudy (anonymous):

\[ y = b^x\] ..............is equivalent to............... (means the exact same thing as) \[\log{b} (y) = x\]

OpenStudy (anonymous):

hmm but it also might want you to just take the log of both sides and simplify for x sorry about being so hazy... just giving my initial ideas :) the only thing I'm super confident about is that you'll need to use logs!

OpenStudy (inkyvoyd):

protip: take the log base 37 of both sides

OpenStudy (inkyvoyd):

you could also take the ln of both sides but that's for losers

OpenStudy (anonymous):

ok take the log of 37 then I would end up with? @inkyvoyd

OpenStudy (inkyvoyd):

x=log_37 12

OpenStudy (inkyvoyd):

then, use a change of base

OpenStudy (inkyvoyd):

\(\Huge \log_a b=\frac{\log_d b}{log_d a}\) you can do anything for d. I usually go with the natural log ln or common log log_10

OpenStudy (anonymous):

So If i were to describe in words I would just say covert this equation to log for and then solve for d? @inkyvoyd

OpenStudy (inkyvoyd):

idk

OpenStudy (inkyvoyd):

someone else answer this, I'm feeling lazy :P

OpenStudy (anonymous):

@inkyvoyd plssssssssss help its getting late and I have to turn this in

OpenStudy (inkyvoyd):

To be honest, I get how to solve for x in here, but I havne't hte slightest idea how to answer this question.

OpenStudy (inkyvoyd):

@Confusionist , plz help plz thx

OpenStudy (anonymous):

@Confusionist plssssssssssss helppppppppp i beggg you

OpenStudy (confusionist):

I'm with you on this one, sir @inkyvoyd. I dont know how to answer this. >_>

OpenStudy (anonymous):

plsssssssssssssssssss @Confusionist

OpenStudy (wolf1728):

Take logs of both sides (I can do this)

OpenStudy (anonymous):

omg omg yes yea

OpenStudy (anonymous):

x =LOG_37 12

OpenStudy (wolf1728):

37^x = 12 log (37^x) = log (12) The log of (37^x) equals x * log (37) So now we have x * log (37) = log (12) x = log (12) / log (37) x = 1.079181246 / 1.5682017241 x = 0.6881648129 and that's that :-)

OpenStudy (anonymous):

Thank you

OpenStudy (wolf1728):

u r welcome - wow lucky I came along huh?